OpenStudy (rayford):

simplify the expression 4/(3x)^-2

OpenStudy (anonymous):

use x^(-a) = 1/x^a

OpenStudy (anonymous):

example x^(-5) = 1/x^5

OpenStudy (jiteshmeghwal9):